# Learning Object Compliance via Young’s Modulus from Single Grasps with Camera-Based Tactile Sensors

### v1

@@ 5.1 Hardware Setup
GelSight Wedge sensors [22] were rigidly attached to a standard parallel robotic gripper with 3D-printed plastic fingers, as depicted in Fig. 5. One sensor is inscribed with markers for tracking surface displacement. Markers help provide information about shear forces and surface displacement, as further explained in Appendix B. By collecting data with different sensors simultaneously, we have the ability to assess the relative necessity of markers in estimating Young’s modulus. Normal force is measured through a gauge embedded in the left finger. Gripper position is directly measured from control software.
@@ 5.2 Data Collection & Training
Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type. Data is fit to functions of the form $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
@@ 6 Experimental Results
We train our neural network with preprocessed data as described in Section 5.2. Initially, we train a model over only rubber objects. This preliminary exercise is performed to evaluate the model’s ability to learn across a constrained dataset. Rubber objects are closest to the Young’s modulus of our sensor. From Eq. 5, we expect it will be easier to discern finer differences in compliance between them. Results over unseen objects for this model are plotted in Fig. 8. The model demonstrates 100% prediction accuracy within an order of magnitude of the ground truth Young’s modulus. The prediction range of this experiment is analogous to previously developed Shore hardness tactile models [16], which are designed to estimate the compliance of soft rubber objects. Shore 00 hardness values evaluated in [16] range from approximately 6 kPa to 2 MPa. Here, we include harder rubber objects as well.
We train our neural network over the entire collected dataset. This learned model is compared to analytical baselines. Predictions across the dataset for each method are plotted in Fig. 9. We observe that as the traditional Hooke’s Law method without tactile sensing fails to distinguish between all material types. Simple elastic and Hertzian analytical methods demonstrate significant improvements, but are still less than 30% accurate. Our learned model showcases dramatic improvements over all other methods.
All estimation methods had more difficulty expressing precise differences in Young’s modulus between harder objects that strayed further from the modulus of the sensor (i.e. plastic, wood, and metal objects). Additionally, we observe decreased overall accuracy from previous evaluation over the rubber subset of our dataset. These results are expected given that harder objects are approximately rigid relative to the soft sensor’s Young’s modulus. The resolution of the tactile sensor is likely not high enough to detect small differences in deformation between this range of compliances.
In Table 1, performance is compared between different estimation methods and inputs. Performance metrics are computed for grasps of both seen and unseen objects. We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset. Predictions are considered sufficiently accurate within the same order of magnitude of ground truth labels. We compute error as the average difference between prediction and ground truth on a log10 scale.
@@
From the ablation study, we observe that our learned model vastly reduces error and improves accuracy over analytical methods. Furthermore, we see that RGB tactile images with markers are favorable for this estimation task. However, we do not observe significant difference in results between models which consider force $F$ and width $w$ measurements. This may be due to the fact that all objects were grasped with nearly identical peak force. Moreover, force is implicitly detectable through tactile images [16]. Importantly, considering analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ is shown to improve learned prediction results. This demonstrates that incorporating analytical modeling improves our ability to discern compliance over a purely learning-based method. Accuracy and error metrics are computed across shape and material categories for each method in Appendix D. We observe that our model performs independent of shape and can generalize well across materials.
@@ 7 Conclusion
We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods. This system can create estimates online from only a single parallel grasp. It has been shown to perform independent of object shape and demonstrates robustness across materials. It can expressively distinguish precise differences in compliance between soft objects. The system is less capable of measuring precise differences between harder, such as plastic and metal, objects which are effectively rigid relative to the soft GelSight sensor.
The estimation system introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth. With more diverse tactile data, the accuracy of the system could be further improved. Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects. As one example, it could be utilized to identify the ripeness of fruits or vegetables [13, 8, 10].
@@ Appendix A Method of Dimensionality Reduction (MDR)
As discussed in Section 3.2, we model contact between an unknown object and our camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory. Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [59]. A mechanical diagram of contact at time $t$ with labeled variables is provided in Fig. 10.
@@ Appendix B Markers
Markers are small black dots painted in an array on the surface of the GelSight gel pad. As the surface of the gel deforms and stretches, these markers move. Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions. Markers have been shown to be effective for measuring shear and normal forces from tactile images [15, 22]. From a broad standpoint, this means we can use markers to increase the amount of dynamic information encoded in a tactile image. Without them, we could not estimate shear force.
In our experiment, we have created one GelSight Wedge [22] with markers painted on the surface. This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus. The difference between tactile images with or without markers for two objects is tabulated in Fig. 11. These markers help track surface displacement, but they may occlude fine features of texture, like screw threads.
@@ Appendix C Processing Grasp Data
Appendix D Dataset Diversity & Performance
The diversity of the collected dataset is characterized through categories of material and shape in Tables 3 and 3. As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset. This is done to bias the dataset toward the Young’s modulus of our sensor’s gel pads. There are significant percentages of metal and plastic objects, making up 25.1% and 24.0% of the dataset respectively. Over 50% of objects in the dataset have irregular shapes. This is done to diversify training data in an attempt and learn Young’s modulus independent of object shape or contact profile. Still, there are substantial percentages of simple shapes for performance evaluation.

### v2

@@ 5.1 Hardware Setup
GelSight Wedge sensors [17] were rigidly attached to a standard parallel robotic gripper with 3D-printed plastic fingers, as depicted in Fig. 5. Although we choose GelSight sensors for our experiment, other flat, camera-based tactile sensors capable of tracking depth in high-resolution would be suitable. One sensor is inscribed with markers. As explained in Appendix B, markers help enhance the dynamic information in a tactile image by tracking surface displacement. Data from only one sensor is required for estimation, but we collect data with both sensors simultaneously to later assess the necessity of markers in estimating Young’s modulus. Additionally, we assume symmetry between fingers in our analytical models. Normal force is measured through a gauge embedded in the left finger. Gripper position is directly measured from control software.
@@ 5.2 Data Collection & Training
Figure 6: Raw force and depth data. Data is collected across grasps for spheres of each material type and fit to $d\sim F^{\frac{2}{3}}$, per our Hertzian model.
@@ 6 Experimental Results
Initially, we train an instance of our model using only rubber objects from our dataset. From Eq. 5, we expect it will be easier to discern fine differences in compliance between rubber objects, as their Young’s modulus is closest to that of our sensor. Estimations across unseen objects for this model are plotted in Fig. 8. The model demonstrates 100% prediction accuracy within an order of magnitude of ground truth Young’s modulus, with an average log10 error of 0.36. We utilize this reduced training configuration to compare our architecture against previously developed Shore hardness models [14]. As detailed in Appendix D, our hybrid system exhibits reduced error over this limited domain.
Subsequently, we train our neural network over the entire collected dataset and compare against analytical baselines. Predictions across unseen objects for each method are plotted in Fig. 9. A traditional Hooke’s Law method without tactile sensing fails to distinguish between material types. Simple elastic and Hertzian analytical models exhibit stronger performance, but remain less than 30% accurate. Our hybrid learned model showcases dramatic improvements over all methods.
In Table 1, performance is compared between different estimation methods and inputs. Performance metrics are computed for grasps of both seen and unseen objects. We evaluate models on a logarithmic scale given the range of Young’s moduli in our dataset. We compute error as the average difference between prediction and ground truth on a log10 scale. Predictions are considered accurate within the same order of magnitude of ground truth labels. This tolerance is deemed sufficient for downstream tasks of produce sorting based on the change in compliance of selected fruits and vegetables during ripening [54, 55, 56, 57].
@@
From the ablation study, we observe that our hybrid model drastically out-performs purely analytical and learning-based methods. RGB tactile images with markers are shown to be most favorable for this estimation task. There is not significant difference in results between models which consider force $F$ and width $w$ measurements. This may be due to the fact that all objects were grasped with identical peak force. Moreover, force is implicitly apparent through tactile difference images [12]. Importantly, analytical estimates $\hat{E}_{\text{elastic}},\hat{E}_{\text{hertz}}$ are shown to improve prediction accuracy by nearly 10%. This indicates that a hybrid approach to compliance estimation is favorable over purely learning-based methods, which have been previously explored [14, 15, 1, 2, 3, 4]. Prediction accuracy across shape and material categories for each method is investigated in Appendix E. We observe that our model performs independent of shape and can generalize well across materials.
@@
Limitations: Our estimation system struggled to express precise differences in Young’s modulus between harder objects that stray further from the modulus of our sensor (i.e. plastic, wood, and metal objects). These results are expected given that harder objects are rigid relative to our soft sensor. The resolution of our tactile sensor is not high enough to detect minute differences in deformation between this range of compliances. Considering this fact, our estimation system exhibits only marginal effective improvement relative to previously developed Shore hardness estimation models [14]. Further work must be done to create a truly universal compliance estimation system.
@@ 7 Conclusion
We have developed a novel system for estimating the Young’s modulus of unknown contacted objects using analytical and data-driven methods. This system can create estimates online from only a single parallel grasp using camera-based tactile sensors. It has been shown to perform independent of object shape and remains robust across material types. It can precisely distinguish between compliances of soft objects, but struggles to expressively differentiate between that of harder objects.
The estimation architecture introduced in this paper could be applied for use with other camera-based tactile sensors capable of tracking depth. With more diverse tactile data, the accuracy of the system could be further improved. Given an increased ability to generalize, this compliance estimation system could be deployed alongside an intelligent high-level controller to inform decision-making in robotic manipulation scenarios by characterizing the compliance of unknown objects. This is particularly valuable when visually similar objects may have distinct material composition. As one example, it could be utilized to identify the ripeness of fruits or vegetables [1, 2, 4].
@@ Appendix A Method of Dimensionality Reduction (MDR)
As discussed in Section 3.2, we model contact between an unknown object and our flat, camera-based tactile sensor as a collision between an elastic sphere and an elastic half-space using Hertzian contact theory. Now, we will walk through the derivation for a constitutive relation using the method of dimensionality reduction (MDR) [58]. A mechanical diagram of contact at time $t$ with labeled variables is provided in Fig. 10.
@@ Appendix B Markers
Markers are small black dots painted in an array on the surface of the GelSight gel pad. As the surface of the gel deforms and stretches, these markers move. Thus, we can track their relative position to map the displacement of the gel pad surface in both shear and normal directions. Markers have been shown to be effective in estimating shear and normal forces from tactile images [12, 17]. However, accurate force estimations require assumptions on contact geometry. In this work, we seek to estimate the compliance of unknown objects. Since we cannot make accurate assumptions on geometry, we measure forces through an external force gauge rather than through our GelSight sensors. In any case, we can use markers to increase the amount of dynamic information encoded in a tactile image. Without them, we could not perceive shear displacement.
In our experiment, we have created one GelSight Wedge [17] with markers painted on the surface. This was done to investigate how beneficial markers may be for estimating a dynamic property of contact like Young’s modulus. The difference between tactile images with or without markers for two objects is displayed in Fig. 11. These markers help track surface displacement, but they may occlude fine features of texture, like screw threads.
@@ Appendix C Processing Grasp Data
Appendix D Comparison Against Shore 00 Hardness Baseline
Previous works have developed learning-based methods to estimate Shore hardness using camera-based tactile sensors without the incorporation of analytical modelling [40, 14, 5, 15, 1]. Notably, Yuan et al. introduced an architecture to estimate Shore 00 hardness of various sample geometries using tactile images captured from a GelSight sensor [14]. To demonstrate the performance of our novel hybrid estimation system, we directly compare it against results from Yuan et al.
As discussed in Section 6, we can constrain our dataset to predict compliance for only rubber objects. There are 82 rubber objects in our dataset of 285 total objects. The compliance of these rubber objects aligns with the Shore 00 hardness scale, which spans approximately from a Young’s modulus of 6 kPa to 2 MPa. We can directly evaluate and compare predictions of Young’s modulus with Shore 00 hardness by utilizing established models that convert between the two measures [53].
Using the rubber subset of our dataset, we replicate experiments conducted by Yuan et al. and compare their quoted results against our own [14]. We use RGB tactile images with markers and analytical estimates as the input to our model. First, we train an instance of our hybrid estimation system to predict compliance using only simple shapes such as spheres, cylinders, and rectangular objects. This model is validated across seen shapes with unseen compliances. Predictions from our model are subsequently converted to Shore 00 hardness to calculate error over the validation set. Second, we train another instance of our hybrid model and validate it across unseen objects with arbitrary or irregular shapes. It is important to note that Yuan et al. evaluate over hand-collected data for the latter experiment, whereas all of our data collection is performed on robot hardware. Results from our experiments are presented in Table 2, where root mean-squared error (RMSE) is reported in units of Shore 00 hardness.
Method Dataset $R^{2}$ RMSE
Shore 00 Hardness Simple Shapes 0.85 10.3
Model [14] Arbitrary Shapes 0.39 18.2
Top-1 Hybrid NN Simple Shapes 0.69 5.9
(Our Method) Arbitrary Shapes 0.37 14.3
Table 2: Shore 00 Hardness Prediction Error Comparison
Our estimation system exhibits decreased error in both experiments across both simple and complex shapes, though the margin of improvement is modest. These results suggest that our novel hybrid estimation architecture offers some performance enhancement over previously developed learning-based approaches for predicting the compliance of soft rubbers. We can likely attribute this decrease in error to the incorporation of analytical modeling.
Appendix E Dataset Diversity & Performance
The diversity of the collected dataset is characterized through categories of material and shape in Tables 4 and 4. As intended, the most common material of objects in the dataset is rubber, which includes 28.6% of the objects in the dataset. This is done to bias the dataset toward the Young’s modulus of our sensor’s gel pads. There are significant percentages of metal and plastic objects, making up 25.1% and 24.0% of the dataset respectively. Over 50% of objects in the dataset have irregular shapes. This is done to diversify training data in an attempt and learn Young’s modulus independent of object shape or contact profile. Still, there are substantial percentages of simple shapes for performance evaluation.